mr.reyes is 20 years older than his son.How old are the father and the son if mr. reyes is 6 times older as his son?
Answer by Cromlix(4381)

mr.reyes is 20 years older than his son
Mr R = 6* Son
R = 6S
mr. reyes is 6 times older as his son
R = 20 + S
Replace R with 6S from first equation.
6S = 20 + S
6S - S = 20
5S = 20
S = 4
Currently,
Father is 24 years old
Son is 4 years old.
Hope this helps.
:-)